A bodyvof mass 12kg travelling at 4.2m/s collide with a sound body of mass 18kg at rest. Calculate their common velocity


1
Expert's answer
2021-10-18T10:57:42-0400

Their common velocity can be found from momentum conservation:


"mv+M*0=(m+M)u,\\\space\\
u=mv/(m+M)=1.68\text{ kg m/s}."